OpenStudy (lgbasallote):

how to integrate \(\int (\sin^{-1} x)^{2} dx\)

OpenStudy (anonymous):

An approach to this is with integral by parts

OpenStudy (australopithecus):

u = arcsin(x) du = 1/(1-x^(2))dx du(1-x^(2)) = dx so now you have (u)^(2)(1-x^(2))

OpenStudy (australopithecus):

sorry \[\int\limits_{}^{}(u)^{2}\sqrt{1-x^{2}}\]

OpenStudy (australopithecus):

it should be u = arcsin(x) du = 1/(1-x^(2))^(1/2)dx du(1-x^(2))^(1/2) = dx

OpenStudy (australopithecus):

Now solve for x and use integration by parts

OpenStudy (australopithecus):

Solve for x u = arcsin(x) sin(u) = sin(arcsin(x)) sin(u) = x

OpenStudy (australopithecus):

Now we have \[\int\limits_{}^{}u^{2}\sqrt{1-\sin^{2}(u)}\] use the identity 1 - sin^(2)(u) = cos^(2) now we have \[\int\limits\limits_{}^{}u^{2}\cos(u)\]

OpenStudy (australopithecus):

Use substitution by parts

OpenStudy (anonymous):

no, you were right..

OpenStudy (australopithecus):

oh yeah :)

OpenStudy (australopithecus):

Thanks

OpenStudy (anonymous):

keep going, I am watching you (:

OpenStudy (australopithecus):

Now use substitution by parts \[\int\limits\limits\limits_{}^{}u^{2}\cos(u)du = u^{2}\sin(u) - \int\limits_{}^{}2usin(u)dx\]

OpenStudy (anonymous):

you mean du

OpenStudy (australopithecus):

yeah ignore my sloppy notation :)

OpenStudy (australopithecus):

use substitution by parts again \[u^{2}\sin(u) + 2ucos(u) + 2\int\limits\limits\limits_{}^{}\cos(u)du\] so u^(2)sin(u) + 2ucos(u) - 2sin(u) + c Now we need to substitute

OpenStudy (anonymous):

u^(2)sin(u)- 2ucos(u) - 2sin(u) + c ?

OpenStudy (australopithecus):

arcsin^(2)(x)x + 2arcsin(x)cos(arcsin(x)) - 2x + c

OpenStudy (lgbasallote):

i think there's a wrong sign there :P

OpenStudy (australopithecus):

the anti-derivative of sin(x) is -cos(x)

OpenStudy (lgbasallote):

the antiderivative of cosx is sinx though haha

OpenStudy (australopithecus):

No I'm pretty sure I'm right it seems right to me

OpenStudy (australopithecus):

\[u^{2}\sin(u) - 2u(-\cos(u)) - 2\int\limits_{}^{} -\cos(u)du\]

OpenStudy (australopithecus):

I really enjoyed that thanks :)

OpenStudy (lgbasallote):

haha lol =))

OpenStudy (australopithecus):

arcsin^(2)(x)x + 2(1-x^(2))^(1/2)arcsin(x) - 2x + c and I was right before with my negative :L cos(u) = (1-sin(u)^(2))^(1/2) = (1-x^(2))^(1/2)

OpenStudy (australopithecus):

the final answer is arcsin^(2)(x)x + 2(1-x^(2))^(1/2)arcsin(x) - 2x + c

OpenStudy (australopithecus):

Going to check with wolfram alpha now :)

OpenStudy (australopithecus):

Confirmed :) https://www.wolframalpha.com/input/?i=integrate+arcsin%28x%29^%282%29 Did you follow sorry it got sketchy in the integration by parts area

OpenStudy (australopithecus):

Hope this was helpful or you just wanted to see someone prove this :)

OpenStudy (lgbasallote):

no..i was really asking someone to solve it hahaha =)))) you just use by parts and variable subs right? no trig sub?

OpenStudy (australopithecus):

well you are kind of doing a trig sub but not really as it is pretty intuitive.

OpenStudy (australopithecus):

remember that cos(u) = \[\sqrt{1-x^{2}}\] sin(u) = x and u = arcsin(x)

OpenStudy (australopithecus):

you need to remember to sub those values into the end equation, I had this same question except it wasn't squared on my calc 1 final, I killed it :)

OpenStudy (australopithecus):

but yes these questions are easy just very long

OpenStudy (australopithecus):

I should really have gone one step at a time instead of skipping steps though

OpenStudy (australopithecus):

sorry if I wasn't any help

OpenStudy (lgbasallote):

nahh..izfine haha